Find a parametrization of the tangent line at the point indicated. r(t) = (1 - 4, 4t, 5t), t = 2
Answer #1

Scle. 8(t) = {1-tt, 4t, 5 t3 » 9t=2 L Lt) = 8lto) + tolto) t=to=2 8(2) = 2-15 ,8 , 40) 8 Lt) = - 4t3, 4, 15t2) 8 (2) = {-40